OpenStudy (surryyy):

Given the graph below, which of the following points would be included in the graph of the inverse of the function? Choose all that apply. (1, 0) (0, 1) (-2, 1) (1, 2)

OpenStudy (anonymous):

01, iinverse ii2 the 2ame a2 rever2e

OpenStudy (anonymous):

2orry iif iim not correct though .-.

imqwerty (imqwerty):

1st-we will write the equation of this graph

imqwerty (imqwerty):

so we choose 2 points which lie on it and then we'll use double point form - \[\frac{ y-y _{1} }{ y _{2 }-y _{1}}=\frac{ x-x _{1} }{ x _{2}-x _{1}}\]where (x1,y1) and (x2,y2) are the 2 points which lie on the graph :)

imqwerty (imqwerty):

so the 2 points which i chose are -(0,1) and (1,-2) substituting the values we get-

imqwerty (imqwerty):

(y-1)/(-2-1)=(x-0)/(1-0) (y-1)/-3 = x y-1=-3x y=1-3x <-----equation of given graph

imqwerty (imqwerty):

to find the equation of the inverse of this graph we just need to switch y and x

imqwerty (imqwerty):

so the equation which we get is- x=1-3y <-----equation of inverse of the given graph

imqwerty (imqwerty):

now u just need to try the options ..........put them in the equation nd if they satisfy then they lie on this inverse graph :)

OpenStudy (surryyy):

So its (1, 0) and (-2, 1)

imqwerty (imqwerty):

yes :)

OpenStudy (surryyy):

Thanks! :)

imqwerty (imqwerty):

no prblem :D